A) at what time t_a will the current in the circuit reach 1/2 of its maximum value?
B) at what time t_b will the energy stored in the inductor reach 1/2 of its maximum value?
Problem 5 (10 POINTS) A battery, ε-30.0V, a resistor, R = 40.00, and an inductor, L-1.5mH, are con- nected in series with an open switch. The switch is then closed at time t = 0. (a) At what time ta will the current in the circuit reach 1/2 of its maximum value? (b) At what time th will the energy stored in the inductor reach 1/2 of its maximum value?
Answer #1

a) The current as a function of time is given by:

I(t) = I_0(1-e^{-Rt/L})

where I_0= \frac{V}{R}

the time ta when the current is 1/2 of its maximum value:

\frac{I_0}{2}= I_0(1-e^{-Rt_a/L})

solving for ta we have:

t_a = \frac{L}{R}l\ln(2) = \frac{1.5\times 10^{-3}}{40.0}\ln(2)=2.599\times 10^{-5}s

b) The time tb when the energy stotred in the inductor reach 1/2 of its maximum value:

U = \frac{1}{2}LI^2

as a function of time:

U(t) = U_0(1-e^{-Rt/L})^2

for tb

\frac{U_0}{2} = U_0(1-e^{-Rt_b/L})^2

\frac{1}{2} = (1-e^{-Rt_b/L})^2

solving for tb we have:

t_b =\frac{L}{R} \ln(\sqrt{2})=\frac{1.5\times 10^{-3}}{40.0} \ln(\sqrt{2})=4.605\times 10^{-5} s